OpenStudy (anonymous):

HELP! Algebra 1 question. Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (anonymous):

Which do you think it is?

OpenStudy (anonymous):

not really sure..

OpenStudy (kidrah69):

y=mx+b

OpenStudy (anonymous):

If two lines are parallel, what is true about those two lines?

OpenStudy (anonymous):

that they never intersect?

OpenStudy (anonymous):

that they have the same slope!

OpenStudy (anonymous):

^ yes what kid rah said!

OpenStudy (kidrah69):

oops

OpenStudy (anonymous):

Hahaha it's oaky

OpenStudy (kidrah69):

\[y-y1=m(x-x1)\]

OpenStudy (anonymous):

You use the point slope formula and remember to plug in the same slope!

OpenStudy (anonymous):

Can you write the equation you came up with? :D

OpenStudy (anonymous):

is the formula kidrah put the point slope formula?

OpenStudy (anonymous):

yep :)

OpenStudy (kidrah69):

\[y-(-6)=1/2(x-4)\]

OpenStudy (anonymous):

Also plug in the points where they intersect

OpenStudy (kidrah69):

now just solve that

OpenStudy (anonymous):

wait ignore that, idk why it put that my connection is freaking out.

OpenStudy (anonymous):

nvm ignore that last statement lol

OpenStudy (anonymous):

what would be the first step to solving that equation?

OpenStudy (kidrah69):

Ok the first step is to use point slop formula that is what i wrote above can you solve that?

OpenStudy (anonymous):

no i meant whats the first step in solving the point slope formula? sorry im not good at math..

OpenStudy (kidrah69):

\[y+6=1/2(x-4)\] -(-6) = 6

OpenStudy (kidrah69):

after that you distribute the 1/2

OpenStudy (anonymous):

wait how did you get -(-6) = 6?

OpenStudy (kidrah69):

Because when you have 2 negatives next to each other it will turn positive

OpenStudy (anonymous):

y+6=(x−2)?

OpenStudy (anonymous):

y+6=1/2(x-2)*

OpenStudy (anonymous):

y+6=(x-2)*

OpenStudy (anonymous):

y+6=1/2x-2*

OpenStudy (kidrah69):

right now combine the like terms

OpenStudy (anonymous):

y+4=1/2x

OpenStudy (kidrah69):

no subtract 6 from both sides we want y=mx+b form

OpenStudy (anonymous):

y=1/2x-8

OpenStudy (kidrah69):

yep :)

OpenStudy (anonymous):

thats it?

OpenStudy (kidrah69):

yep

OpenStudy (anonymous):

oh ok lol, thanks alot man appreciate it!

OpenStudy (kidrah69):

lol *girl :P and welcome to openstudy! :)

OpenStudy (anonymous):

oh my bad lol, and thanks!